OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 5 ?

OpenStudy (anonymous):

what is \[\log\bigg(\frac{A}{B}\bigg) \] equal to?

OpenStudy (anonymous):

I don;t know.

OpenStudy (anonymous):

Ah, Remember that logs are essentially exponents and follow their math rules. Just like \[\frac{x^a}{x^n}=x^{a-n}\] \[log\bigg( \frac{A}{B}\bigg) = log(A) - log(B)\] and \[log\bigg(AB\bigg) = log(A) + log(B)\]